OpenStudy (anonymous):

PLEASE HELP Domain of g(x) = square root of (x-1/x+7) Type your answer in interval notation. Simplify your answer. Use integers of fractions for numbers in the expression.

OpenStudy (campbell_st):

\[\sqrt{\frac{x - a}{x + 7}}\] first thing to consider is that x cannot equal 7, as this will give a zero denominator..

OpenStudy (anonymous):

I meant x-1 not x-a

OpenStudy (campbell_st):

ok... so x can't be -7 and the who fraction needs to be positive as you can't take the square root of a negative number so x- 1 >= 0 so the domain would be x >= 1

OpenStudy (anonymous):

okay thanks!